OpenStudy (anonymous):

the length of the batter's box on a softball field is 1 ft less than twice the width. the area of the batter's box is 15 ft^2. Find the length and width of the batter's box. Can the formula for a rectangle apply to this?

OpenStudy (anonymous):

A=L*W L= 2W-1 15 = (2W-1)*W 15= 2W^2-W 0= 2W^2 -W-15 (2W+5)(W-3) W=-5/2 or W=3 the width can not negative than W=3 is solution L= (2*3)-1 L= 5 check: A=3*5 A=15 click good answer if I deserve medal

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

thank you!